OpenStudy (anonymous):

find the roots of x^2+12x+38=0

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

ROOTS: \[x = -6-i \sqrt(2)\] \[x = -6+i \sqrt(2)\]

OpenStudy (anonymous):

USE QUADRATIC FORMULA OR COMPLETING THE SQUARE

OpenStudy (anonymous):